# Non-hydrodynamic response in QCD-like plasma

### 13 Dec 2023 v1

@@ 1 Introduction
Another motivation of our work is to further investigate and elaborate on the notion of the extended hydrodynamic regime (EHR) proposed by us in Ref. Ke:2022tqf , This scenario corresponds to the situation that a high-frequency sound exists that governs the medium’s response outside the hydrodynamic regime on the one hand, and its dispersion is not described by naively extending conventional hydrodynamic theories outside the hydrodynamic scale on the other hand. In Ref. Ke:2022tqf , we demonstrate the presence of this scenario for strongly coupled ${\cal N}=4$ super-symmetric Yang-Mills (SYM) theory and RTA kinetic theory. We also show that the sound dispersion can be captured by a ”deformed” version of MIS, namely MIS*, with a suitable choice of model parameters. In this work, we show detailed analysis indicating that the same MIS* equations can describe not only dispersion relation but also response function outside the hydrodynamic regime.
@@
This paper is organized as follows. In sec. 2.1, we derive a general relation between the response function describing the response induced by initial disturbance and the retarded correlator describing the medium’s reaction to the metric perturbation, see eq. (10) (or eq. (11)). While it directly follows from the standard linear response theory, it seems new in the literature. We then compute the energy-momentum response function in a static background from RTA kinetic theory and hydrodynamic theories in sec. 2.2 and 2.3 respective. In sec. 3.1, we analyze the excitations that dominate the response for the RTA kinetic theory. In doing so, we review the concept of EHR and discuss its generality. Then we review the construction of MIS* in sec. 3.2. Sec. 3.3 compares kinetic response and that from different hydrodynamic theories. In sec. 4, we set up the formalism to calculate similar response functions in a Bjorken expanding background and present results in 5. The sensitivity of the Euclidean correlator to non-hydrodynamic response is discussed and illustrated in sec. 6. Finally, we summarize in sec. 7.
@@ 2.2 Kinetic theory
$\delta f_{\textrm{eq}}=e^{-p^{0}/T_{0}}\left(-p_{0}\delta\beta+\beta_{0}\,{\bm{p}}\cdot\delta{\bm{u}}\right)\,$ (21)
@@
where we have introduced the short-handed notation $f_{0}\equiv e^{\beta_{0}p_{0}}$ for the background equilibrium distribution and $\bm{v}=\hat{p}$ is the single particle velocity. We shall solve eq. (22) under the initial condition that the distribution is in local equilibrium:
@@
$\delta f(\omega,\bm{k};\bm{v})=\frac{i\,\beta_{0}p_{0}f_{0}\tau_{R}^{-1}}{(\omega-\bm{v}\cdot\bar{\bm{k}}+i\tau_{R}^{-1})}\,\left[\left(-\frac{\delta\beta(\omega,\bm{k})}{\beta_{0}}+\delta\bm{u}(\omega,\bm{k})\cdot\bm{v}\right)+\frac{\delta f_{I}(\bm{k};\bm{v})}{\beta_{0}p_{0}f_{0}\tau_{R}^{-1}}\right]\,.$ (24)
@@ 3.1 Excitations
Figure 2: RTA response function in the time domain (solid curves) at three different values of $\bar{k}=k\tau_{R}$. They are compared with the contribution from the collective modes $\omega_{L}(k),\omega_{T}(k)$ (dashed curves) which are conventional sound and shear modes in hydrodynamic regime (see eq. (57) and the text). In the extended hydrodynamic regime for RTA kinetic theory where $\bar{k}\geq 1$, those modes are referred to as high-frequency sound and shear mode. (Left): the longitudinal channel where we have rescaled the time by $\Omega=|{\rm Re}\Omega_{L}(k)|$ (Right): the transverse channel where we have rescaled the time by $\Omega=|{\rm Im}\Omega_{T}(k)|$.
@@
In Fig. 2 which exhibits the RTA response function vs $\Delta t$ for three values of $\bar{k}$ ($\bar{k}<\bar{k}_{O}$) in the longitudinal and transverse channel, we have shown the contribution from the poles and branch cuts separately. Remarkably, the collective modes have a prevalent impact on the behavior of response functions even at the non-hydrodynamic gradient, especially at sufficiently large $t$. To understand why this happens, we compared the damping rate of RTA collective modes with the relaxation rate of quasi-particle excitations shown in Fig. 3. We refer to the difference between the two, $\Delta\Gamma(k)=\tau_{R}^{-1}-{\rm Im}(\omega(k))$, as the gap. If the gap is greater than zero, the contribution from quasi-particle excitations will be suppressed by $\exp(-\Delta\Gamma\Delta t)$, eventually making high-frequency hydrodynamic modes dominate the response. The gap remains open unless the value of $\bar{k}$ is close to $\bar{k}_{O}$, as seen in Fig. 3, which explains the importance of high-frequency sound and shear modes outside the hydrodynamic regime.
@@ 3.1.1 Extended hydrodynamic regime: generalities
Furthermore, high-frequency sound mode is not uncommon in strong-coupled gauge theories that can be calculated using holography duality. For example, in strongly coupled ${\cal N}=4$ super-symmetric Yang-Mills theory in large $N_{c}$ limit (SYM), which captures many features of QCD in strong coupling, the sound mode exists at any finite gradient and has a smaller than damping rate than other excitations (quasi-normal modes) Amado:2008ji ; Fuini:2016qsc . This similarity between kinetic theory and Strongly coupled SYM theory is remarkable, given that the coupling strength is drastically different between the two.
@@ 3.2 The construction of $*{}^{*}$
where $\eta_{1,2},\tau_{1,2}$ are model parameters. Without losing generality, we shall assume that $\tau_{1}<\tau_{2}$, meaning $\pi_{2}$ evolves faster than $\pi_{1}$. In the long time limit, $\pi^{\mu\nu}$ should approach to its value in the first-order hydrodynamics (41), yielding the following constraint
@@
In Fig. 4, we show the longitudinal (the upper panel) and transverse (the lower panel) response function vs $k$ for $\Delta t=2\,\tau_{R}$ (left), $\Delta t=8\,\tau_{R}$ (middle) and $\Delta t=16\,\tau_{R}$ (right). They represent early-time, intermediate-time, and late-time responses, respectively. For kinetic theory, ${\cal G}_{L}$ exhibits the damped oscillation behavior while ${\cal G}_{T}$ is purely dissipative. As $k$ increases, Hydrodynamic and RTA kinetic theory start to differentiate from each other. MIS theory, which includes higher-order gradient effects, only modestly improves the description when $k$ is small. Outside the hydrodynamic regime, both first-order and MIS theory are inadequate to describe the RTA response (see also Ref. Hong:2010at ) The discrepancy can be primarily explained by the difference in the dispersion of collective modes among those theories. For instance, the phase velocity $v(k)$ of RTA sound mode rises from the conformal sound velocity $\sqrt{1/3}\approx 0.57$ up to $0.8$ as $k$ increases, and the sound attenuation rate also rises with $k$. In contrast, (extrapolated) first-order hydrodynamics does not give a supersonic phase velocity, and the sound damping rate grows faster than the RTA theory. Consequently, as shown in Fig. 4, first-order hydrodynamics is less oscillatory with a smaller amplitude than the RTA response function. As for the MIS theory, the phase velocity of the sound mode also increases, but its attenuation rate will be saturated instead of growing as $k$ becomes larger. This explains the oscillating behavior in the MIS response function at large $k$ where RTA sound mode has already been damped out.
@@
In Fig. 4, we illustrate the comparison between RTA and MIS* response function with a representative combination of model parameters $(\gamma,\delta)=(0.25,0.8)$. Despite the model’s simplicity, the agreement in the sound channel is rather remarkable. In the shear channel, the agreement is less impressive. Pursuing improvement in this channel is interesting, which we leave to future work. That said, we stress that MIS* is still very useful in its current form when the hydrodynamization happens first in the shear channel since, in this case, capturing the EHR response in the sound channel is more important. This is what has been accomplished by MIS*.
@@ 4.1 Response function for Bjorken background
where we assume the background is under Bjorken expanding. Accordingly, background EMT $T^{\mu\nu}_{0}(\tau)$ only depends on the Bjorken proper time $\tau=\sqrt{t^{2}+z^{2}}$ and takes form
@@ 4.2 Kinetic theory calculation
In this section, we present the calculation of the response function for a Bjorken expanding plasma using RTA kinetic equation (16). In Bjorken coordinate and for the slice at middle spatial rapidity $\eta_{s}=0$, the kinetic equation (16) reads
@@ 4.3 Hydrodynamic theory calculation
$\partial_{\tau}\pi_{0}=-\frac{1}{\tau_{\pi}}\,(\pi_{0}-\frac{2\eta}{3\tau})-(\frac{4}{3}+\frac{\tilde{\lambda}_{1}}{3})\,\frac{\pi_{0}}{\tau}$ (94)
@@ 5.3 Response in Fourier space
Figure 6: The energy-momentum response function $G_{\epsilon\epsilon}$, $G_{L\epsilon}$, $G_{TT}$, defined in eqs. (71) and (72) (c.f. their analogs in the static background in Fig. 6), are plotted in the solid black curves from the first to the third row as a function of $k\Delta\tau$. Those functions are rescaled by multiplying $(\tau/\tau^{\prime})^{4/3},(\tau/\tau^{\prime}),(\tau/\tau^{\prime})$ respectively, see text. The results of RTA kinetic theory, first-order hydrodynamics, and MIS theory are shown in the solid, dotted, and dashed curves, respectively. The MIS* results are plotted in the red band, computed from $0.80.2<\delta<0.3,0.7<\gamma<$ 0.8. Three values of $\Delta\tau$ represent the response function at an early, intermediate, and late time (from left to right). To test relation (101), we plot $(\tau/\tau^{\prime})G_{LL}$ and $(\tau/\tau^{\prime})G_{\epsilon L}/3$ from the kinetic theory in the black dashed curve in the first and second row, respectively. They are to be compared with rescaled $G_{\epsilon\epsilon}$ and $G_{L\epsilon}$.
@@
$(\frac{\tau}{\tau_{0}})^{4/3}G_{\epsilon\epsilon}\approx(\frac{\tau}{\tau_{0}})G_{LL}\qquad G_{L\epsilon}\approx(\frac{1}{3})G_{\epsilon L}\,.$ (101)
@@ 5.4 Response in real space
Finally, we note that ${\cal G}_{LL}(r)$, defined in relation (4.1) and shown in the second row of Fig. 7, receive contributions from both sound and shear modes, but it exhibits the sound peak in $r/\Delta\tau$ as its main feature. We have also checked that this peak is also in ${\cal G}_{TT}$. This further makes us believe the describing non-hydrodynamic sound propagation is more important than that for the shear mode.
@@ 6 Non-hydrodynamic response from the Euclidean correlators
where $\tau_{E}$ is the Euclidean time. There have been extensive efforts in extracting the low-frequency behavior of spectral density from the lattice QCD calculation of $G_{E}$ to determine transport coefficients such as shear viscosity non-perturbatively (see Ref. Meyer:2011gj for a review). Nevertheless, since the characteristic frequency of the kernel that convoluted with the spectral density in eq. (102) is of the order $\omega_{T}=\pi T$, the resulting Euclidean correlators are notoriously insensitive to the features in hydrodynamic regimes where the typical gradient and frequency are expected to be smaller than $\omega_{T}$
@@
$k_{{\rm EHR}}=(1\text{--}2)\,(\frac{\eta}{s})^{-1}\,c^{2}_{s}\,T\,.$ (103)
@@
Here, $\rho_{{\rm EHR}}$ can obtained by computing $G_{R}$ from the MIS* response function through the relation (10). We find
@@ 7 Summary and outlook
We have also compared the energy-momentum response function at static and Bjorken-expanding. We learned that at a non-hydrodynamic gradient, the response as a function of proper time is not very sensitive to the bulk background profile. This indicates that one may use the response functions obtained in the static background to estimate that in the expanding bulk profile, allowing for the simplification in the numerical studies. We also learned that although hydrodynamics would eventually describe those response functions, the ”hydrodynamization time” of the response function would be numerically large. This in turn suggests that jet-medium observable might be employed to extract the medium’s properties outside the hydrodynamic regime, and MIS* theory, or its extension, can be applied within this context.

### 09 Apr 2024 v2

@@ 1 Introduction
Another motivation of our work is to further investigate and elaborate on the notion of the extended hydrodynamic regime (EHR) proposed by us in Ref. Ke:2022tqf , This scenario corresponds to the situation that a high-frequency sound exists that governs the medium’s response outside the hydrodynamic regime on the one hand, and its dispersion is not described by naively extending conventional hydrodynamic theories outside the hydrodynamic scale on the other hand. In Ref. Ke:2022tqf , we demonstrate the presence of this scenario for strongly coupled ${\cal N}=4$ super-symmetric Yang-Mills (SYM) theory and RTA kinetic theory. We also show that the sound dispersion can be captured by a “deformed” version of MIS, namely MIS*, with a suitable choice of model parameters. In this work, we show detailed analysis indicating that the same MIS* equations can describe not only dispersion relation but also response function outside the hydrodynamic regime.
@@
This paper is organized as follows. In sec. 2.1, we derive a general relation between the response function describing the response induced by initial disturbance and the retarded correlator describing the medium’s reaction to the metric perturbation, see eq. (10) (or eq. (11)). While it directly follows from the standard linear response theory, it seems new in the literature. We then compute the energy-momentum response function in a static background from RTA kinetic theory and hydrodynamic theories in sec. 2.2 and 2.3 respectively. In sec. 3.1, we analyze the excitations that dominate the response for the RTA kinetic theory. In doing so, we review the concept of EHR and discuss its generality. Then we review the construction of MIS* in sec. 3.2. Sec. 3.3 compares kinetic response and that from different hydrodynamic theories. In sec. 4, we set up the formalism to calculate similar response functions in a Bjorken expanding background and present results in 5. The sensitivity of the Euclidean correlator to non-hydrodynamic response is discussed and illustrated in sec. 6. Finally, we summarize in sec. 7.
@@ 2.2 Kinetic theory
$\delta f_{\textrm{eq}}=e^{-\beta_{0}p^{0}}\left(-p_{0}\delta\beta+\beta_{0}\,{\bm{p}}\cdot\delta{\bm{u}}\right)\,$ (21)
@@
where we have introduced the short-handed notation $f_{0}\equiv e^{-\beta_{0}p^{0}}$ for the background equilibrium distribution and $\bm{v}=\hat{p}$ is the single particle velocity. We shall solve eq. (22) under the initial condition that the distribution is in local equilibrium:
@@
$\delta f(\bar{\omega},\bar{k};\bm{v})=\frac{i\,\beta_{0}p_{0}f_{0}\tau_{R}^{-1}}{(\omega-\bm{v}\cdot\bar{\bm{k}}+i\tau_{R}^{-1})}\,\left[\left(-\frac{\delta\beta(\bar{\omega},\bar{\bm{k}})}{\beta_{0}}+\delta\bm{u}(\bar{\omega},\bar{\bm{k}})\cdot\bm{v}\right)+\frac{\delta f_{I}(\bar{\bm{k}};\bm{v})}{\beta_{0}p_{0}f_{0}\tau_{R}^{-1}}\right]\,.$ (24)
@@ 3.1 Excitations
Figure 2: RTA response function in the time domain (solid curves) at three different values of $\bar{k}=k\tau_{R}$. They are compared with the contribution from the collective modes $\omega_{L}(k),\omega_{T}(k)$ (dashed curves), which are conventional sound and shear modes in hydrodynamic regime (see eq. (57) and the text). In the extended hydrodynamic regime for RTA kinetic theory where $\bar{k}\geq 1$, those modes are referred to as high-frequency sound and shear mode. (Left): the longitudinal channel where we have rescaled the time by $\Omega=|{\rm Re}\Omega_{L}(k)|$ (Right): the transverse channel where we have rescaled the time by $\Omega=|{\rm Im}\Omega_{T}(k)|$.
@@
In Fig. 2, which exhibits the RTA response function vs $\Delta t$ for three values of $\bar{k}$ ($\bar{k}<\bar{k}_{O}$) in the longitudinal and transverse channel, we have shown the contribution from the poles and branch cuts separately. Remarkably, the collective modes have a prevalent impact on the behavior of response functions even at the non-hydrodynamic gradient, especially at sufficiently large $t$. To understand why this happens, we compared the damping rate of RTA collective modes with the relaxation rate of quasi-particle excitations shown in Fig. 3. We refer to the difference between the two, $\Delta\Gamma(k)=\tau_{R}^{-1}-{\rm Im}(\omega(k))$, as the gap. If the gap is greater than zero, the contribution from quasi-particle excitations will be suppressed by $\exp(-\Delta\Gamma\Delta t)$, eventually making high-frequency hydrodynamic modes dominate the response. The gap remains open unless the value of $\bar{k}$ is close to $\bar{k}_{O}$, as seen in Fig. 3, which explains the importance of high-frequency sound and shear modes outside the hydrodynamic regime.
@@ 3.1.1 Extended hydrodynamic regime: generalities
Furthermore, high-frequency sound mode is not uncommon in strongly coupled gauge theories that can be calculated using holography duality. For example, in the strongly coupled ${\cal N}=4$ super-symmetric Yang-Mills theory in large $N_{c}$ limit (SYM), which captures many features of QCD in strong coupling, the sound mode exists at any finite gradient and has a smaller than damping rate than other excitations (quasi-normal modes) Amado:2008ji ; Fuini:2016qsc . This similarity between kinetic theory and strongly coupled SYM theory is remarkable, given that the coupling strength is drastically different between the two.
@@ 3.2 The construction of $*{}^{*}$
where $\eta_{1,2},\tau_{1,2}$ are model parameters. Without losing generality, we shall assume that $\tau_{1}<\tau_{2}$, meaning $\pi_{2}$ relaxes faster than $\pi_{1}$. In the long time limit, $\pi^{\mu\nu}$ should approach to its value in the first-order hydrodynamics (41), yielding the following constraint
@@
In Fig. 4, we show the longitudinal (the upper panel) and transverse (the lower panel) response function vs $k$ for $\Delta t=2\,\tau_{R}$ (left), $\Delta t=8\,\tau_{R}$ (middle) and $\Delta t=16\,\tau_{R}$ (right). They represent early-time, intermediate-time, and late-time responses, respectively. For kinetic theory, ${\cal G}_{L}$ exhibits the damped oscillation behavior while ${\cal G}_{T}$ is purely dissipative. As $k$ increases, Hydrodynamic and RTA kinetic theory start to differentiate from each other. MIS theory, which includes higher-order gradient effects, only modestly improves the description when $k$ is small. Outside the hydrodynamic regime, both first-order and MIS theory are inadequate to describe the RTA response (see also Ref. Hong:2010at ). The discrepancy can be primarily explained by the difference in the dispersion of collective modes among those theories. For instance, the phase velocity $v(k)$ of RTA sound mode rises from the conformal sound velocity $\sqrt{1/3}\approx 0.57$ up to $0.8$ as $k$ increases, and the sound attenuation rate also rises with $k$. In contrast, (extrapolated) first-order hydrodynamics does not give a supersonic phase velocity, and the sound damping rate grows faster than the RTA theory. Consequently, as shown in Fig. 4, first-order hydrodynamics is less oscillatory with a smaller amplitude than the RTA response function. As for the MIS theory, the phase velocity of the sound mode also increases, but its attenuation rate will be saturated instead of growing as $k$ becomes larger. This explains the oscillating behavior in the MIS response function at large $k$ where RTA sound mode has already been damped out.
@@
In Fig. 4, we illustrate the comparison between RTA and MIS* response function with a representative combination of model parameters $(\gamma,\delta)=(0.25,0.8)$. Despite the model’s simplicity, the agreement in the sound channel is rather remarkable. In the shear channel, the agreement is less impressive. Pursuing improvement in this channel is interesting, which we leave to future work. That said, we stress that MIS* is still very useful in its current form when hydrodynamization happens first in the shear channel, since capturing the EHR response in the sound channel is more important in this case. This is what has been accomplished by MIS*.
@@ 4.1 Response function for Bjorken background
where we assume the background is under Bjorken expanding. Accordingly, background EMT $T^{\mu\nu}_{0}(\tau)$ only depends on the Bjorken proper time $\tau=\sqrt{t^{2}-z^{2}}$ and takes form
@@ 4.2 Kinetic theory calculation
In this section, we present the calculation of the response function for a Bjorken expanding plasma using RTA kinetic equation (16). In Bjorken coordinate the kinetic equation (16) reads
@@ 4.3 Hydrodynamic theory calculation
$\partial_{\tau}\pi_{0}=-\frac{1}{\tau_{\pi}}\,\left(\pi_{0}-\frac{2\eta}{3\tau}\right)-\left(\frac{4}{3}+\frac{\tilde{\lambda}_{1}}{3}\right)\,\frac{\pi_{0}}{\tau}$ (94)
@@ 5.3 Response in Fourier space
Figure 6: The energy-momentum response function $G_{\epsilon\epsilon}$, $G_{L\epsilon}$, $G_{TT}$, defined in eqs. (71) and (72) (c.f. their analogs in the static background in Fig. 6), are plotted in the solid black curves from the first to the third row as a function of $k\Delta\tau$. Those functions are rescaled by multiplying $(\tau/\tau^{\prime})^{4/3},(\tau/\tau^{\prime}),(\tau/\tau^{\prime})$ respectively, see text. The results of RTA kinetic theory, first-order hydrodynamics, and MIS theory are shown in the solid, dotted, and dashed curves, respectively. The MIS* results are plotted as red bands, computed from $0.80.2<\delta<0.3,0.7<\gamma<$ 0.8. Three values of $\Delta\tau$ represent the response function at an early, intermediate, and late time (from left to right). To test relation (101), we plot $(\tau/\tau^{\prime})G_{LL}$ and $(\tau/\tau^{\prime})G_{\epsilon L}/3$ from the kinetic theory in the black dashed curve in the first and second row, respectively. They are to be compared with rescaled $G_{\epsilon\epsilon}$ and $G_{L\epsilon}$.
@@
$\left(\frac{\tau}{\tau_{0}}\right)^{4/3}G_{\epsilon\epsilon}\approx\frac{\tau}{\tau_{0}}G_{LL}\qquad G_{L\epsilon}\approx\frac{1}{3}G_{\epsilon L}\,.$ (101)
@@ 5.4 Response in real space
Finally, we note that ${\cal G}_{LL}(r)$, defined in relation (4.1) and shown in the second row of Fig. 7, receive contributions from both sound and shear modes, but it exhibits the sound peak in $r/\Delta\tau$ as its main feature. We have also checked that this peak is also in ${\cal G}_{TT}$. This further makes us believe that describing non-hydrodynamic sound propagation is more important than that for the shear mode.
@@ 6 Non-hydrodynamic response from the Euclidean correlators
where $\tau_{E}$ is the Euclidean time. There have been extensive efforts in extracting the low-frequency behavior of spectral density from the lattice QCD calculation of $G_{E}$ to determine transport coefficients such as shear viscosity non-perturbatively (see Ref. Meyer:2011gj for a review). Nevertheless, since the characteristic frequency of the kernel that is convoluted with the spectral density in eq. (102) is of the order $\omega_{T}=\pi T$, the resulting Euclidean correlators are notoriously insensitive to the features in hydrodynamic regimes where the typical gradient and frequency are expected to be smaller than $\omega_{T}$.
@@
$k_{{\rm EHR}}=(1\text{--}2)\,\left(\frac{\eta}{s}\right)^{-1}\,c^{2}_{s}\,T\,.$ (103)
@@
Here, $\rho_{{\rm EHR}}$ can be obtained by computing $G_{R}$ from the MIS* response function through the relation (10). We find
@@ 7 Summary and outlook
We have also compared the energy-momentum response function at static and Bjorken-expanding. We learned that at a non-hydrodynamic gradient, the response as a function of proper time is not very sensitive to the bulk background profile. This indicates that one may use the response functions obtained in the static background to estimate that in the expanding bulk profile, allowing for the simplification in the numerical studies. We also learned that although hydrodynamics would eventually describe those response functions, the ”hydrodynamization time” of the response function would be numerically large. This, in turn, suggests that jet-medium observables might be employed to extract the medium’s properties outside the hydrodynamic regime, and MIS* theory, or its extension, can be applied within this context.